15. Assume TRO) 150+50Q 502 and TC(Q) 100 6Q Create a table in Excel showing TR(Q), MR(Q) , TC(Q), MC(Q), π(Q), and marginal profit, as Q goes from 0 to 10 in increments of1 b. Identify the level of Q that maximizes profit. c. At the level of Q that maximizes profit, what is the relation between MR and MC? d. At the level of Q that maximizes profit, what is marginal profit? e. Check your answer in part b using calculus
Answer #1

a)

Q TR TC MR = ∆TR/∆Q MC= ∆TC/∆Q Profit = TR - TC Marginal Profit = ∆Profit/∆Q
0 150 100 - - 50 -
1 195 106 45 6 89 39
2 230 112 35 6 118 29
3 255 118 25 6 137 19
4 270 124 15 6 146 9
5 275 130 5 6 145 -1
6 270 136 -5 6 134 -11
7 255 142 -15 6 113 -21
8 230 148 -25 6 82 -31
9 195 154 -35 6 41 -41
10 150 160 -45 6 -10 -51

b) Profit is maximized at the output level of 4.

c) At the profit maximizing level of output, MR > MC.

d) At the profit maximizing level of output, Marginal Profit is $9.

e) Profit = TR - TC = (150 + 50Q - 5Q2) - (100 + 6Q) = = 150 + 50Q - 5Q2 - 100 - 6Q = 50 + 44Q - 5Q2

Marginal Profit = 44 - 10Q

Profit is maximized When Marginal Profit i > or = 0

44 - 10Q > or = 0

44 > or = 10Q

44 / 10 > or = Q

4.4 > = Q

Thus, if we round it, we can find Q = 4.